your bank balance is $987.27 .If you write a check to buy a pair of shoes ,your balance would be $967.63.How much do the shoes cost? Write a linear equation that models the problem.
Answer by checkley71(8403)

987.27-S=967.63 or -S=967.63-987.27 or -S=-19.64 or SHOES=$19.64